OpenStudy (anonymous):

Logarithmic equation .3 = x + 10^(-14)*e^(x/.026)

OpenStudy (zzr0ck3r):

\(.3=x+\frac{e^\frac{x}{.026}}{10^{14}}\)?

OpenStudy (anonymous):

\[.3 = x + 10^{-14}e ^{\frac{ x }{ .026 }}\]

OpenStudy (anonymous):

yah

OpenStudy (zzr0ck3r):

hmm, not sure if you can solve this.

OpenStudy (anonymous):

yah not sure how to do it because if you take the log of both sides then the lone x gets messed up and vice versa

OpenStudy (zzr0ck3r):

the answer is way above my pay grade. https://www.wolframalpha.com/input/?i=.3+%3D+x+%2B+10^%28-14%29*e^%28x%2F.026%29 scroll down to where it says solution

OpenStudy (anonymous):

the answer is .3 but not sure how they got it

OpenStudy (zzr0ck3r):

thats an approximate, they used the link I sent you.

OpenStudy (zzr0ck3r):

there is no basic analytic way to solve this, you need to be pretty high up there is math to solve it.

OpenStudy (anonymous):

take x on LHs and solve it . you will get something like \[\log(.3-x) = y \] now you will sovle it by taking antilog

OpenStudy (anonymous):

lol yah it seems whatever is on the left hand side is the answer as long as it is within the domain

OpenStudy (zzr0ck3r):

its like solving this 3^x = 2x

OpenStudy (zzr0ck3r):

@crazysingh that wont work, keep going and you will see.

OpenStudy (anonymous):

wait let me solve than @zzr0ck3r

OpenStudy (zzr0ck3r):

ok:)

OpenStudy (anonymous):

lol

OpenStudy (ranga):

10^(-14) is a very very small number close to zero and it can be ignored leaving you with x = 0.3

OpenStudy (anonymous):

@ranga yah that is what i'm thinking

OpenStudy (anonymous):

this is for an electrical circuits class so that is usually what they do with this kind of stuff anyhow

OpenStudy (anonymous):

if the number is small enough then ignore

OpenStudy (anonymous):

ok i think i'm just going to go with that. thank you veruy much everyone

OpenStudy (ranga):

You are welcome.

OpenStudy (zzr0ck3r):

\(x\approx0.002(150-13*W_n\frac{e^{\frac{150}{13}}}{2600000000000})\) where W is the analytic continuation of the product log function and n is natural

OpenStudy (zzr0ck3r):

i have no idea what that means:)

OpenStudy (anonymous):

yah that is way over my head

OpenStudy (anonymous):

lol